OpenStudy (anonymous):

The Burlington Phone Company charges 8 cents per minute for any phone call within the continental US. Its rival, the Try Our Phone Instead Company, charges only 4 cents per minute for any phone call within the continental US, but it also charges a 28-cent “per call charge” every time you make a call. Write a system of equations representing these facts. Calculate how long a phone call must be to make the phone charges of both companies identical. If your average phone call is 11 minutes long, which company should you choose? Make sure you show the basis for your decision.

OpenStudy (anonymous):

Can anyone help me with this?

OpenStudy (gw2011):

First, I am making an assuption that this for one call. Let B = Burlington Phone Company Let T = Try Our Phone Let CH = The Charge Let M = Minutes Let C = Per call charge B: CH = 0.08M T: Ch = 0.04M + 0.28C These are your equations Identical phone charges: 0.08M = 0.04M + 0.28C 0.08M - 0.04M = 0.28C 0.04M = (0.28) (1) 0.04M = 0.28 M = 0.28/0.04 M = 7 minutes 11 Minute Phone Call: B: (0.08) (11) = 0.88 T: (0.04) (11) + (0.28) (1) = 0.72 Since Try Our Phone will cost 72 cents and Burlington will cost 88 cents, it is cheaper to use Try Ou Phone. At least these are the answers I got based upon the information given above.

OpenStudy (anonymous):

ty soooooo much